### v1

# Theorem-Carrying-Transaction: Runtime Certification to Ensure Safety for Smart Contract Transactions

Abstract
Security bugs and trapdoors in smart contracts have been impacting the Ethereum community since its inception. Conceptually, the 1.45-million Ethereum’s contracts form a single “gigantic program” whose behaviors are determined by the complex reference-topology between the contracts. Can the Ethereum community be assured that this gigantic program conforms to its design-level safety properties, despite unforeseeable code-level intricacies? Static code verification is inadequate due to the program’s gigantic scale and high polymorphism. In this paper, we present a viable technological roadmap for the community toward this ambitious goal. Our technology, called Theorem-Carrying-Transaction (TCT), combines the benefits of concrete execution and symbolic proofs. Under the TCT protocol, every transaction carries a theorem that proves its adherence to the specified properties in the invoked contracts, and the runtime system checks the theorem before executing the transaction. Once a property is specified in a contract, it can be treated confidently as an unconditional guarantee made by the contract. As case studies, we demonstrate that TCT secures token contracts without foreseeing code-level intricacies like integer overflow and reentrancy. TCT is also successfully applied to a Uniswap codebase, showcasing a complex decentralized finance (DeFi) scenario. Our prototype incurs a negligible runtime overhead, two orders of magnitude lower than a state-of-the-art approach.
1 Introduction
Decentralized computing technologies have great potential to transform our society. However, a significant obstacle is the lack of robust assurance for smart contracts (i.e., application code executed on a blockchain platform). Numerous high-profile incidents have resulted in substantial financial losses due to unintentional vulnerabilities and deliberate logic trapdoors in smart contracts. These losses are irreversible due to the principle of code-is-law in decentralized computing [1]. For instance, the 2016 DAO attack led to a loss of 3.6 million ethers and a permanent hard fork of the Ethereum network [2]. The 2018 BeautyChain attack basically wiped out the entire asset base of the BeautyChain ecosystem [3]. Two attacks in 2022 exploited lending contracts to over-borrow ethers worth 11.7 million USD [4].
Vision. The community has developed techniques to detect code patterns leading to vulnerabilities, such as the integer-overflow and reentrancy patterns exploited by the BeautyChain and DAO attacks [2][3]. However, a more fundamental approach to trustworthiness is to ensure that an implementation conforms to its design-level safety properties, rather than to detect known patterns. Imagine traveling back to 2016 when the two patterns are still unknown. Suppose the designers of the ERC20 token standard [5] (which specifies the interface for fungible-token currencies) want to ensure that all ERC20 token contracts are secure against these unforeseen attacks, what can they do? It turns out that the property in Prop. 1 is sufficient to thwart both the DAO and BeautyChain attacks. This property asserts that every account balance (i.e., balances[x]) of this token contract should be non-negative, and the sum of all balances should equal the total supply of this currency (i.e., totalSupply). It simply states the designers’ intention, as opposed to enumerating all non-intentions as code-level anti-patterns. Section 3 will explain how this straightforward property defeats the unforeseen attacks.
[…]
Design-level properties are much more powerful than simple comparisons of concrete values (e.g.“x==y”). They may contain quantifiers, set/map operations, relations between encrypted variables, code-reflection properties (e.g., “no variable other than x can be modified”), etc. We envision a future in which contract designers and developers are empowered to express these powerful properties as the “higher law” to govern the contract code, and the platform ensures these properties so that the contracts are impervious to potential code bugs or trapdoors. We believe that a platform providing such high assurance will substantially bolster the public’s confidence on smart contract technology.
Problem formulation. Smart contracts represent a unique programming paradigm. Conceptually, the whole community is co-developing a single gigantic “program”, as in Figure 1: everyone, even an attacker, can be a programmer and/or a user of this program; every contract is an OO object with a cryptographically strong identity called the “contract address”, i.e. an object reference. Any interaction between two objects is made by calling the target object via a reference. The platform (e.g. Ethereum VM, or EVM) sequentially (i.e. no concurrency) executes transactions, each of which is triggered by a user-invocation of a public method in an object. In 2022, the number of contracts on Ethereum’s mainnet had reached 1.45 million [6]. It is worth emphasizing that they are not 1.45 million individual programs, but a single program with 1.45 million objects interconnected!
Figure 1: Smart contracts on Ethereum form a single gigantic “program” with about 1.45-million OO objects.
For brevity, we refer to this gigantic program as $GP$. $GP$ has been self-evolving since 2015 and is highly polymorphic: (1) a transaction’s semantics is determined by the reference topology between the objects; (2) a transaction can create new contracts and references. The reference topology is a result of real-world business and trust relations, not implied by the static code. Therefore, static code alone is insufficient for reasoning about $GP$’s behaviors.
We use $P$ to denote the set of design-level properties co-defined by the whole community, which can be tens/hundreds of thousands of properties like Prop. 1. Hence, our technological goal can be formulated as: ensuring $GP$’s conformance to $P$. This formulation is very different from how a smart contract verification problem is typically formulated. Usually, a verification technique considers “the program” to be one or a few “contracts of interest”, denoted as $C$. $C$ is examined within a hypothetical environment, which is an enormous simplification of the huge $(GP-C)$ code in our formulation. It also assumes some hypothetical malicious users to invoke $C$ repeatedly. The verification serves as a helpful advisory to developers because it flags potential vulnerabilities, but the results are w.r.t. these vastly simplified assumptions. In contrast, our goal is a strong assurance – all properties in $P$ are indeed held true in reality, so everyone can trust them as strong contractual promises.
Our technology. This is obviously an ambitious goal. Static verification is inadequate because of the gigantic scale and high polymorphism of $GP$. In this paper, we present a runtime certification approach and a roadmap for the Ethereum community to deploy it. The technology is named Theorem-Carrying-Transaction (TCT), inspired by Proof-Carrying-Authentication [7] and the concolic testing methodology [8]. In TCT, every transaction must carry a theorem that proves its adherence to the properties in every contract to be executed in the transaction. The runtime platform checks the validity of the theorem before allowing the transaction to be actually executed. The concolic nature of TCT means that it performs a symbolic verification on the concrete straight-line code trace of a transaction, which makes the corresponding theorem efficiently provable. Because all variables are symbolic, a proven theorem has the generality to cover all future transactions that follow the same code trace. This enables TCT to achieve near-zero amortized runtime overhead via a theorem reuse mechanism.
We will show that TCT achieves behavioral subtyping [9], a decades-long aspiration for secure OO programming. It mandates the properties of a base class to hold in all descendant classes. For example, if the designers specify Prop. 1 for the base class ERC20, they can rest assured that all future contracts derived from it will satisfy the property. Behavioral subtyping also gives the implementers a safety-certification mechanism. The fact that a contract is derived from a base class is essentially a certificate to prove to everyone that it meets the specified safety requirements.
TCT’s trusted computing base (TCB) is entirely at the infrastructure level, i.e., the VM, the compiler, and the code verifier. It makes no assumption at the contract-code level, e.g., about the (static) code, the (dynamic) object-reference topology, or any party who claims that a contract has been formally verified. Of course, TCT shares the same caveat as all verification techniques – safety is only as high as the safety spec goes. We hope TCT motivates the community to shift focus from examining code-level intricacies to defining the design-level safety spec.
Prototype and highlights of results. We have built a prototype to demonstrate the TCT approach. In this prototype, a programmer uses the NatSpec notation of Solidity [10] to specify contract invariants and function post-conditions. TCT’s core module proves that these invariants and post-conditions will be satisfied under a given hypothesis (i.e., condition) w.r.t. a concrete EVM code trace (a.k.a. code path). This component consists of the techniques to obtain the code trace, to symbolically execute it and to convert the result into a straight-line code in a verification-amenable language (i.e., Boogie language [11] in our prototype). The hypothesis checking is performed on the Ethereum node by modifying the Go-Ethereum implementation (Geth) [12].
We demonstrate TCT using an ERC20 token contract that can be exploited by the integer overflow and reentrancy patterns. With intuitive hypotheses specified about the contract’s entry functions, our technology automatically proves the desired invariants, which thwart the attacks. To show its generality, the same technology is successfully applied to the transfer transactions of the top-5 real-world ERC20 tokens. Moreover, we showcase Uniswap-V2 as a highly complex real-world scenario. To the best of our knowledge, this is the first code-level verification for Uniswap’s essential properties – inverse proportionality for swap operations and proportionality for liquidity operations. Each transaction executes about 7000 $\sim$ 9000 EVM instructions across five different contracts. Indeed, we expose subtle conditions under which the properties fail to hold. They may lead to vulnerabilities if an investment strategy depends on Uniswap’s claimed properties.
The runtime overhead of TCT is negligible. For the token contract and the Uniswap-V2 system, the overhead measurements are only 0.20% and 0.57% on Geth 1.13.0, respectively. The former is two orders of magnitude lower than a state-of-the-art approach, which inserts runtime checks into token contracts via code instrumentation [13].
2 Background
We briefly describe the techniques that inspire TCT.
PCC and PCA. Proof-Carrying Code (PCC) [14] is designed to allow a trusted party (e.g.an OS kernel) to safely accept a code module from an untrusted producer (e.g. a third-party driver). The novelty is that the producer bears the responsibility to prove the code’s conformance to the safety properties declared by the trusted party, who only needs to check the proof. Proof-Carrying Authentication (PCA) [7] is an authentication mechanism inspired by PCC. Real-world authentication policies, such as those involving “certificate authority”, “delegate”, the “speaksfor” relation, etc., are essentially statements about statements, so the authentication logic is higher-order, which is undecidable. PCA requires every request to carry a proof to convince the server that it is permitted by the authentication logic, so the server only needs to perform a decidable proof-checking task. PCA emphasizes its logic extensibility. Anyone can submit a proof of any theorem, e.g. “Charlie is a certificate authority”. Once the proof is confirmed, the theorem is accepted into the logic as a lemma, which can be used by the proofs of future theorems. Between PCC and PCA, the TCT concept is closer to the latter – a theorem attests to the safety of every runtime request rather than a code module.
Concolic testing. Another inspiration is from concolic testing, short for “concrete symbolic testing”, is a software testing technique that uses symbolic reasoning to analyze concrete code paths. It aims to systematically explore a program’s execution paths to increase code coverage and identify potential bugs. Concolic testing is particularly useful for finding complex and hard-to-detect software vulnerabilities, making it a valuable tool in software verification and security. For example, SAGE [15] is a concolic fuzz-testing technology that has found numerous bugs in Microsoft Windows and Office executables. The demonstrated scalability is inspiring for us because the code complexity of smart contracts is much lower than real-world binary executables.
3 Motivating Example
To motivate our work, this section describes a token contract in Solidity that combines two vulnerabilities – the first is exploitable via integer overflow, similar to the one described in reference [16]; the second is exploitable via reentrancy, similar to the DAO attack [2]. In Figure 2, the middle portion shows the contract code MultiVulnToken. Smart contract execution is sequential and transactional. The require statement checks a condition: if it is false, the execution is reverted so that the EVM state remains the same as if the execution did not occur.
Figure 2: Attacks via integer overflow and reentrancy.
3.1 Vulnerabilities and attacks
In Attack 1, the attacker issues a transaction by calling transferProxy (in transition A). Unexpectedly, arguments _value and _fee are set to $2^{255}$+1 and $2^{255}$. In the line require(balances[_from] >= _fee + _value), the two values are added, which yields result 1 due to integer overflow, so the require statement is satisfied, i.e., the execution is not reverted and the rest of the function also follows through. Address msg.sender is the attacker’s address, and address _to is also specified by the attacker. When the function returns, the balances of the two addresses become enormous (as transition B). This causes extreme inflation, so the attacker gets almost the entire asset of this token ecosystem.
Attack 2 results in the attacker receiving 50 tokens in one of his accounts by clearing his other account that has only 5 tokens. It is achieved by making reentrant calls 10 times. The steps are as follows. First, the attacker creates a contract to call function clear in MultiVulnToken. Inside the function, msg.sender refers to the attacker’s contract. Therefore, the function call msg.sender.receiveNotification (transition C) calls into the attacker’s contract, who calls function clear again (transition D). This unexpected code sequence results in balances[_to] += bal executed 10 times.
Consider from the perspective of a static code verifier. The behavior of clear is not fully defined because it calls into another contract using an address (i.emsg.sender). The verifier can perhaps err by conservatively flagging all such calls as “potentially unsafe”. However, calling contract via an address reference is very common in smart contracts. Imagine another call businessPartner.foo(), where businessPartner is a variable only writable by the contract owner. Treating it as an arbitrary address, thus flagging the code as “potentially unsafe”, is problematic. With its concolic nature, TCT is not subject to this dillema.
3.2 Theorem to prove safety of a transaction
The notion of theorem is important in TCT. We explain it using the code paths in the two attacks. For Attack 1, the code path of the transaction is shown in 1, using the Boogie language [11]. A require check in Solidity corresponds to an assume in Boogie, because passing a require check means that the code path is subject to the constraint from the symbolic reasoning perspective. The operators + and - are modeled as functions evmadd and evmsub, correspond to opcodes ADD and SUB in the Ethereum yellow paper [17].
tmp1 := evmadd(_fee, _value);
assume balances[_from] >= tmp1;
tmp2 := evmadd(balances[_to], _value);
assume tmp2 >= balances[_to];
tmp3 := evmadd(balances[msg.sender], _fee);
assume tmp3 >= balances[msg.sender];
balances[_to] := evmadd(balances[_to], _value);
balances[msg.sender] := evmadd(balances[msg.sender], _fee);
balances[_from] := evmsub(balances[_from], tmp1);
Code 1: Code path of Attack 1 in the Boogie language.
2 lists some axioms in TCT’s basic library. They are not specific to the discussed case. In the axioms, m is a map from address to int. Notation m[a:=v] denotes a map that is almost identical to m, but its index a contains the value v.
function sum(m: [address] int) returns (int);
axiom forall m: [address] int, a:address ,v:int::
sum(m[a:=v]) == sum(m) - m[a] + v;
axiom forall m: [address] int ::
((forall a:address :: 0 <= m[a]) ==>
(forall a:address :: m[a] <= sum(m)));
function evmadd(a,b:int) returns (int);
axiom forall a,b:uint256 ::
a+b < 2^256 && a+b >= 0 ==> evmadd(a,b) == a+b;
axiom forall a,b: uint256 ::
a+b >= 2^256 ==> evmadd(a,b) == a+b-2^256;
function sub(a,b:int) returns (int);
axiom forall a,b: uint256 ::
a-b < 2^256 && a >= b ==> evmsub(a,b) == a-b;
axiom forall a,b: uint256 ::
a < b ==> evmsub(a,b) == a-b+2^256;
Code 2: Axioms of mathematical operations.
Because Prop. 1 is specified as invariants of ERC20, the derived MultiVulnToken contract should obey them. Invariants are conditions every execution should maintain, i.e., assuming the invariants are true in the beginning, they should be true in the end. 3 represents the proof obligation of the code path in 1. The first two inv lines assume the invariants, and the last two inv lines assert them. The hyp line (i.e., hypothesis) specifies a condition $\varphi(v,s)$, where $v$ represents parameters of the transaction and $s$ represents the blockchain storage states. The hypothesis is the condition under which the invariants will hold.
hyp: assume $\varphi$(v,s);
inv: assume sum(balances) == totalSupply;
inv: assume forall x:address :: 0 <= balances[x] && balances[x] <= totalSupply;
tmp1 := evmadd(_fee, _value);
assume balances[_from] >= tmp1;
tmp2 := evmadd(balances[_to], _value);
assume tmp2 >= balances[_to];
tmp3 := evmadd(balances[msg.sender], _fee);
assume tmp3 >= balances[msg.sender];
balances[_to] := evmadd(balances[_to], _value);
balances[msg.sender] := evmadd(balances[msg.sender], _fee);
balances[_from] := evmsub(balances[_from], tmp1);
inv: assert sum(balances) == totalSupply;
inv: assert forall x:address :: 0 <= balances[x] && balances[x] <= totalSupply;
Code 3: Proof obligation of the code path in 1.
Suppose $\varphi(v,s)$ is (0 $\leq$ totalSupply < $\texttt{2}^{\texttt{255}}$) $\land$ (0 $\leq$ _value < $\texttt{2}^{\texttt{255}}$) $\land$ (0 $\leq$ _fee < $\texttt{2}^{\texttt{255}}$), which consists of the contract variable totalSupply in $s$ and the parameters _value and _fee in $v$ of the entry function transferProxy. With the hypothesis, Boogie can successfully verify 3, meaning that the theorem $\tau$ in 4 is valid.
$\tau$ := ( f = multiVulnToken::transferProxy,
$\varphi$(v,s) = (0 $\leq$ totalSupply < $\texttt{2}^{\texttt{255}}$) $\land$
(0 $\leq$ _value < $\texttt{2}^{\texttt{255}}$) $\land$ (0 $\leq$ _fee < $\texttt{2}^{\texttt{255}}$),
ph = 0xf222eb3c699342dcf8f4f64a86ba0862e
29fc8d5d28e21940ee0b32cf3eb06ca
Code 4: An example theorem for transferProxy.
In the theorem $\tau$, the first part $f$ specifies the entry function of the transaction that the theorem pertains to. For readability of this paper, we use the name multiVulnToken::transferProxy, but in reality, the contract is identified by its address, and the function is identified by a 4-byte hexical value called “function selector” [18]. The second part is the hypothesis $\varphi(v,s)$. The third part is the path hash, which is a Keccak-256 hash identifying an EVM code trace. Section 5.4 will describe how it is computed in our current prototype. In summary, the general form of a theorem is $\tau:=(f,\varphi(v,s),ph)$, which means: for every transaction invoked by calling $f$ when $\varphi$ is satisfied, if it is completed (i.e., not reverted by EVM) and its path hash equals $ph$, then all the assertions along the code path are guaranteed to hold. Obviously, this theorem is not applicable to the transaction in Attack 1 because it violates (0 $\leq$ _value < $\texttt{2}^{\texttt{255}}$) $\land$ (0 $\leq$ _fee < $\texttt{2}^{\texttt{255}}$).
The generality of the theorem can be understood from three aspects. First, the hypothesis not only specifies the conditions for the three variables, but also implicitly states that nothing else matters, such as other parameters in the call or other storage variables. For example, addresses msg.sender, _from and _to can be arbitrarily aliased to each other, and the assertions will still hold. The theorem is thus a powerful statement. Second, the generality is far enough for any instance of the token contract in reality, as it is hard to imagine why a cryptocurrency needs a $2^{255}$ total supply (equivalent to $7*10^{66}$ per person in the world). Unlike static code verifiers, which hinge on the ability to expose unexpected boundary conditions, TCT enables programmers to focus on conditions only within the boundary of their intentions. Third, we will explain that TCT utilizes generality of the hypotheses to reuse theorems. It is crucial for TCT’s near-zero runtime overhead.
hyp: assume $\varphi$(v,s);
inv: assume sum(balances) == totalSupply;
inv: assume forall x:address :: 0 <= balances[x] && balances[x] <= totalSupply;
bal := balances[msg.sender];
* tmp1 := evmadd(balances[_to], bal);
* balances[_to] := tmp1;
/* Start of reentrancy */
tmp2 := evmadd(balances[_to], bal);
balances[_to] := tmp2;
--- repeat ---
tmp10 := evmadd(balances[_to], bal);
balances[_to] := tmp10;
/* End of reentrancy */
balances[msg.sender] := 0;
inv: assert sum(balances) == totalSupply;
inv: assert forall x:address :: 0 <= balances[x] && balances[x] <= totalSupply;
Code 5: Proof obligation corresponding to Attack 2.
Theorem about the reentrant code path. We now do a similar exercise about a transaction entering via clear(...). If it does not exploit the reentrancy pattern (i.e. if transition D in Figure 2 becomes a return), then a hypothesis like $\varphi(v,s)$ := (0  $\leq$  totalSupply < $\texttt{2}^{\texttt{255}}$) $\land$ (_to $\neq$ msg.sender) is sufficient to ensure the invariants. However, if the transaction exploits the reentrancy pattern as in Attack 2, the code path is different, as shown in 5. The two lines marked with “*” are repeated 9 more times. Hence the theorem for the non-reentrancy transaction is not appliable to the reentrancy one, because of the path hash mismatch. The programmer is unlikely to anticipate the reentrancy code path before the year 2016, but it is fine because security of TCT does not require the programmer to anticipate every non-intention. Since it is the attacker who discovers this unintended code path, it is the attacker’s obligation to prove the corresponding theorem. The aforementioned $\varphi(v,s)$ will fail the proof. The attacker can make the hypothesis $\varphi(v,s)$ := (balances[msg.sender] = 0) to prove the theorem, but it makes the attack harmless as the transaction is basically a no-op repeated 10 times.
4 The TCT Protocol
The TCT protocol requires only small extensions from the current Ethereum system. The protocol consists of three parties as shown in Figure 3 – the transaction issuer (e.g., MetaMask [19]), a Web-API service (e.g., infura.io) and the Ethereum network (i.e., the blockchain). In today’s Ethereum ecosystem, the Web-API service provides various functionalities, including gas-estimate for a transaction and failure warnings. It consists of a Web-API interface and a node connecting to the Ethereum network. In TCT, the Web-API service provides additional functionalities for storing and retrieving theorems. The most unique component for TCT is the theorem repository on each node, which contains the theorems that have been proven.
Trusted computing base. TCT’s trusted computing base (TCB) is entirely at the infrastructure level. It does not include any contract-level assumption (e.g. about the reference-topology, contracts of interest, contracts that have been formally verified, etc), because TCT faithfully treats the entire GP as the target program. Our TCB includes Ethereum’s TCB and TCT-additional TCB. The former includes the Ethereum protocol, EVM, and the verified-compilation to attest a contract’s EVM code to its claimed source code [20]. The latter includes the TCT protocol (in this section), the Boogie verifier and a module that translates a trace from EVM to Boogie (in Section 5).
Protocol flow A. The first protocol flow is shown in Figure 3, which describes how the transaction issuer submits a transaction TX to the Ethereum network. First, TX is submitted to the Web-API service. In addition to the existing functionalities (e.g., gas-estimation, etc.), the service tries to find a theorem applicable to TX. Applicability means that the parameters ($v$) of TX and the storage states ($s$) satisfy the hypothesis $\varphi(v,s)$ of the theorem, and the path hash of executing TX equals the one in the theorem. Suppose such a theorem $\tau$ is found, the Web-API service sends $\tau$ to the transaction issuer, who then submits the tuple (TX, $\tau$) to the blockchain. Every blockchain node ensures the existence of $\tau$ in the repository, checks $\varphi(v,s)$, then executes TX. When the execution is completed, the actual path hash is compared against the one in theorem $\tau$. If they are not equal, the execution is reverted.
Like the current Ethereum protocol, the TCT protocol does not strictly require the interaction with the Web-API service before submitting a transaction. If the issuer already knows $\tau$, the protocol flow can start from “Send(TX,$\tau$)”.
Figure 3: Protocol flow A – when a transaction has an applicable theorem to prove its safety.
Protocol flow B. The flow in Figure 4 is used for any person, likely a programmer, to add a theorem about a known code trace into the theorem repository. CT represents the EVM bytecode sequence of the code trace. The issuer submits tuple (CT, $\tau$) to the blockchain. Each node computes the path hash of CT, and checks it against the one in the theorem. If they do not match, the theorem is rejected. Otherwise, the protocol proceeds. Section 5 will describe our technology that converts CT and $\tau$ into a Boogie code trace. The Boogie verifier is used to prove the code trace. If successful, the theorem is accepted into the repository. Note that the verification step burns a certain amount of gas to prevent denial-of-service attacks.
The goal of protocol flow B is for a programmer to pre-populate the repo with all anticipated theorems when the contract is deployed. The code traces can be recorded when the contract is tested on a local EVM or other test environments.
Figure 4: Protocol flow B – adding a theorem about a code trace into the repository.
Protocol flow C. In protocol flow A, an applicable theorem is found for TX. If no such theorem is found, the steps are shown as protocol flow C in Figure 5, which combines flow A and flow B. When the node receives $(TX,\tau)$, it test-runs TX to get the EVM code trace CT, then follows the same steps in flow B to prove and accept $\tau$ into the theorem repo. The remaining steps are the same as in flow A, which execute TX for real.
Figure 5: Protocol flow C – when a transaction has no applicable theorem to prove its safety.
Expectation about the three flows. We expect that flow A is the primary flow to process real users’ transactions, because most transactions are expected to be normal and have applicable theorems in the repository. Most theorems are accepted via flow B before the contracts process real user’s transactions, because the programmers are supposed to run test cases that cover all normal scenarios they can imagine. Flow C happens only when a real user issues a transaction that goes through a code path unseen in all test cases, which is likely an attack, e.g., the reentrancy attack in Section 3. However, if the transaction is legitimate, the user needs to take the burden to provide a new theorem, which should be very rare for a well-tested contract. Even in this case, the only real burden for the user is to provide the hypothesis $\varphi(v,s)$ for the theorem. Proving a straight-line code trace is easy. Our experience with Boogie is that all traces we have encountered are proven automatically.
Optimizations for low storage and communication overheads. To store theorems efficiently in the repository, all theorems that share the same entry function and hypothesis but differ in their path hashes can be combined as a multi-path theorem, which states that, if the hypothesis is satisfied, then any code path in this multi-path set is safe. With this design, the number of theorems is the number of hypotheses, regardless of the number of possible code paths under each hypothesis. The amortized cost to store a new code path is only one Keccak-256 value, i.e., 32 bytes.
The instructions in a code path do not need to be stored, because they are only used during the theorem-proving time. Once it is done, the protocol no longer needs the exact EVM instructions, so the path hash suffices. Once a theorem is stored in the repository, a Keccak-256 hash can be computed for it. We call it theorem hash. In all network messages, $\tau$ can be represented by its theorem hash. Because most transactions go through flow A, we can conclude that the network overhead of TCT is 64 bytes per transaction, due to the messages “($\tau$)” and “Send(TX,$\tau$)”.
5 Prototype
In this section, we describe a prototype of TCT that has implemented all the essential mechanisms. The prototype is tested using real-world ERC20 contracts and a sophisticated application – Uniswap-V2, which we will describe in Section 6. These case-studies demonstrate that TCT is applicable to complex real-world multi-contract transactions, proving design-level properties that are difficult for traditional verification approaches.
5.1 Specifying properties in Solidity
The properties are specified using the NatSpec notation of Solidity [10]. A contract invariant is specified using the tag @custom:invariant above the contract’s definition, and a function postcondition is specified using the tag @custom:postcondition above the function’s definition. For example, 6 specifies the invariants for the ERC20 contract. The invariants could also be specified for the MultiVulnToken contract, but we specify them for ERC20 to show that properties of a base class are ensured by TCT for all derived classes, which demonstrates behavioral subtyping.
///@custom:invariant forall x:address :: (0 <= this.balances[x] && this.balances[x] <= this.totalSupply)
///@custom:invariant sum(this.balances) == this.totalSupply
contract ERC20 {
address public owner;
uint256 public totalSupply
mapping(address => uint256) balances;
Function declarations such as transfer(...), allowance(...), transferFrom(...), ...
Code 6: Specifying the invariants for ERC20.
5.2 Proving a theorem for a transaction
Figure 6 shows how a theorem about a transaction is proven in the prototype system. We use Remix IDE [21] as the transaction issuer. The node consists of a Geth executable and other components. From Remix, we deploy a contract and issue a transaction. Geth returns a TX-hash that identifies the transaction. Note that TX-hash is an Ethereum concept, not to be confused with our TCT concept path hash. Based on the debug_traceTransaction feature of Geth, we build a trace fetcher, which uses TX-hash to get the EVM code trace.
Figure 6: Proving a theorem for a concrete code path.
The main component we build is the concolic executor, which will be described next. The execution result of it is a straight-line Boogie code like the one in 3. The input includes the EVM code trace, the theorem, and Solidity compilation symbol information. The properties specified in the NatSpec notation are in the symbol information. When TCT is deployed, the symbol information should come from a verified-compilation service, such as the Etherscan service named Verify & Publish Contract Source Code [20]. The service attests that a contract’s EVM code is compiled from the claimed source code by an official Solidity compiler.
5.3 Concolic executor
The concolic executor is basically an EVM, which sequentially executes the instructions in a code trace. The main difference from a classical EVM is that the computation of each instruction is performed symbolically. The data structure that we define to hold a symbolic value is named symbolic value tree, or SVT.
EVM is a stack-based machine: operands are pushed to the top of the stack, a computation is performed, the operands are popped, then the result is pushed back in. Figure 7 shows an ADD instruction, assuming arguments named a and b holding concrete values 1 and 2, respectively. The left side shows the concrete computation 1+2=3. On the right side, each box with round corners is an SVT node. The stack does not hold concrete values, but references to SVT nodes. In this case, the top element after the operation points to the tree with the root EVMADD and children a and b. For readability, we denote the tree as EVMADD(a, b).
Figure 7: Concrete execution vs. symbolic execution.
Example. The concolic executor requires substantial understanding about the Solidity specification [18] and the EVM specification [17]. Due to space constraints, we skip many details and use the EVM code segment in Figure 8 as an example. The figure shows the initial stack and the stack before 925 JUMPI. Essentially, the code pushes two elements onto the stack top: the first is 0x03a3 (i.e., decimal value 931); the second is ISZERO (LT (SLOAD (MapElement (MultiVulnToken.balances, Partial32B((12, 31), _from))), EVMADD(_fee, _value))).
Figure 8: Execution of an EVM code segment.
A thorough understanding of the EVM code is not required for this section, but some clarifications may help.
  • In the code, [ff]*20 denotes twenty bytes of 0xff, which is used to mask a 32-byte value representing a 20-byte Ethereum address (see instructions 882-904). The SVT subtree is Partial32B((12,31), _from), denoting the lowest 20 bytes of the 32-byte value _from.
  • The EVM storage is a key-value pair. Both the key and the value are 32-byte. The key is obtained by computing a SHA3 over the slot number of the map variable and the index of the element according to reference [22]. In our example, the slot number is 2 (i.e., 878 PUSH1 02), which corresponds to the map variable balances based on the compilation symbol info of MultiVulnToken. The index is Partial32B((12,31), _from). Hence, the MapElement subtree is equivalent to balances[_from] at the Solidity level.
  • SLOAD loads a value from the EVM storage using a key. LT is the less-then comparison. ISZERO(LT(x,y)) is equivalent to $\neg(x<y)$.
Boogie code generation. When the concolic executor gets to 925 JUMPI, which is a branch, it calls the code generation functionality (a.k.a. CodeGen). The fact that the next instruction is 931 (i.e., the first stack element 0x03a3) indicates that the branch is taken, so the branch condition (i.e., the second stack element) should be assumed true. CodeGen performs a post-order traversal of the SVT to generate the following Boogie code. Note that Boogie does not use the object-oriented (OO) notation “obj.var” to represent a variable in an object. Instead, it is modeled as a dictionary item ClassName.var[objRef], as described in reference [11]. Thus, the first line would mean entry_contract.balances[_from] if an OO notation was used. With 7, we can see that the EVM code segment in Figure 8 is equivalent to the Solidity statement require (balances[_from] >= _fee + _value) in Figure 2.
tmp1 := MultiVulnToken.balances[entry_contract][_from];
tmp2 := evmadd(_fee,_value);
tmp3 := (tmp1 < tmp2);
tmp4 := !tmp3;
assume tmp4;
Code 7: The Boogie code generated from Figure 8.
Inter-contract calls. A significant advantage of TCT’s concolic approach is the easiness to handle a transaction involving multiple contracts. The reentrancy transaction in Section 3 is a simple example. In Section 6, we will show that a Uniswap transaction usually involves five contracts. These contracts call each other through address-references, so it is challenging for a static code analyzer to resolve the addresses of all potential callees. TCT does not face this challenge, as every call is resolved in the concrete trace.
We implement three components to support inter-contract calls. First, we give every contract its own stack and memory like in a classical EVM. Second, we follow the Ethereum yellow paper [17] to implement call instructions (e.g., CALL, STATICCALL) and return instructions (e.g., RETURN, STOP). This requires correctly modeling EVM’s marshalling and unmarshalling mechanisms for call arguments. Third, the call stack also needs to hold symbolic values, because a callee contract address is often a result of the caller contract’s computation. Like the data stack in Figure 8, we implement the call stack so that its elements are also SVT references.
The entire multi-contract transaction is generated as a single Boogie function, in which every contract variable is fully qualified by its contract. For example, Multi VulnToken.balances[entry_contract][_from] is the Boogie representation of balances[_from] of the contract at the bottom of the call stack, i.e., entry_contract.
5.4 Concrete checking by Geth
We have described how theorem proving is prototyped in TCT. Whether a theorem is applicable to a transaction requires concretely checking hypothesis $\varphi(v,s)$ and computing a path hash $ph$. We prototype these in Geth 1.13.0.
Hypothesis-checking code is hooked into the beginning of function applyTransaction in the Geth core. We follow the Solidity documentation [22] to get the layout of variables in a contract. We use the Geth function StateDB::GetState to read contract variables.
To compute the path hash, we add a byte buffer to the Interpreter struct in the EVM module. We then modify the file instructions.go so that a branch instruction (i.e., JUMPI) records the target program counter, and a call instruction (i.e., CALL, STATICCALL, DELEGATECALL, CALLCODE) records the callee address. The path hash, which is the Keccak-256 hash of the buffer, is computed at the end of function applyTransaction.
As TCT’s name suggests, every transaction carries a theorem. This is implemented by adding a theorem_hash field, described in the last paragraph in Section 4, in the Transaction struct in the Geth core.
6 Scaling up from ERC20 to Uniswap
ERC20 token is a major application on Ethereum. Section 3 shows one example. To show TCT’s generality, we have applied it to the transfer transactions of the top-5 ERC20 contracts according to Etherscan [23], namely USDT, BNB, USDC, stETH and TONCOIN. These transfers are not more complex than what we show in Section 3. The observation supports our premise – although smart contract’s static code can be complex, each transaction is usually short and simple for symbolic verification, thus TCT can scale to sophisticated cross-contract scenarios.
To test our approach on scenarios that are significantly more complex, we apply TCT to a codebase of the Uniswap decentralized exchange. Uniswap is the most extensively used smart contract application on Ethereum based on the amount of gas burnt [24] (as of May 2024). We now present a very concise explanation about the design of Uniswap’s basic mechanism, based on which we explain how TCT is applied.
6.1 Concise explanation of Uniswap’s design
Uniswap’s core design is an automatic market maker (AMM) algorithm to allow users holding different ERC20 tokens to form a market, which dynamically and automatically sets prices for these tokens. Assuming most users want to maximize their gains from the trades, the prices will reach an equilibrium that represents the fair market prices.
Figure 9: The design of Uniswap’s AMM algorithm.
Uniswap has evolved to version V3 [25], with version V4 in development [26]. We study version V2 [27] because it already contains the essentials in the AMM mechanism, whereas V3’s focus is to add a notion of “concentrated liquidity” to augment the AMM, and V4’s focus is to make Uniswap customizable by developers.
There are two types of users: liquidity providers (LPs) and traders. LPs create pairs (a.k.a. pools in V3) as reserves to facilitate trading of two different ERC20 tokens. Figure 9 shows a pair for TokenA and TokenB reserves. Suppose, at the current moment, the pair holds $x_{1}$ amount of TokenA and $y_{1}$ amount of TokenB in reserve. This is denoted as a black dot at $(x_{1},y_{1})$. LPs can add or remove amounts of TokenA and TokenB to and from a created pair. These operations are add liquidity and remove liquidity. The algorithm requires the amounts to be changed proportionally, as shown by the dashed line $x_{2}/x_{1}=y_{2}/y_{1}$. When an LP adds liquidity to the pair, the system mints an amount of liquidity token, not shown in Figure 9, as the reward for contributing to this pair. The liquidity token is also an ERC20 token. The amount is proportional to the total liquidity contribution. Thus, if its amounts before and after are $z_{1}$ and $z_{2}$, respectively, then the proportionality $x_{2}/x_{1}=y_{2}/y_{1}=z_{2}/z_{1}$ holds. For the reverse direction, when an LP removes liquidity, the system burns the specified amount of liquidity token and transfers the proportional amounts of TokenA and TokenB from the pair back to the LP. Again, $x_{2}/x_{1}=y_{2}/y_{1}=z_{2}/z_{1}$ holds.
The other user type is trader. Traders do not swap tokens with each other, but with the corresponding pair. For example, if traders want to swap TokenA for TokenB or vice versa, they invoke the swap operation with the TokenA-TokenB pair in Figure 9. The AMM algorithm requires the reserve amounts to maintain inverse proportionality, which is also known as the constant-product relation. Suppose the current point is $(x_{1},y_{1})$. A swap operation can move the point to $(x^{\prime}_{1},y^{\prime}_{1})$ along the hyperbola $x_{1}\times y_{1}=x^{\prime}_{1}\times y^{\prime}_{1}$. From the trader’s perspective, they give $\Delta x_{1}=x^{\prime}_{1}-x_{1}$ amount of TokenA to the pair and gets $\Delta y_{1}=y_{1}-y^{\prime}_{1}$ amount of TokenB from the pair. The relative price is $\Delta x_{1}/\Delta y_{1}$. If traders try to profit from trading, the pair’s reserves (i.e., the black dot) will settle in an equilibrium representing the market’s consensus on the relative price between the two tokens. When the market is in action, the operations of LPs and traders interleave, so the dot can move to any point inside the quadrant, e.g., along the path $(x_{1},y_{1})\rightarrow(x_{2},y_{2})\rightarrow(x^{\prime}_{2},y^{\prime}_{2})$.
Safety and fairness. The proportionality and inverse proportionality are the foundation of Uniswap’s design. The founders of Uniswap contracted a company to formally analyze the design and published the analysis report [28, 29]. LP’s proportionality ensures that every LP can get the proportional share of the reserves back at any time. Trader’s inverse proportionality ensures that, if a trader swaps $\Delta x$ TokenA for $\Delta y$ TokenB, anybody can immediately swap $\Delta y$ TokenB to get $\Delta x$ TokenA. We believe that these are the best to say about the algorithm being “safe” and “fair”. For a currency exchange market, we cannot demand “safety” and “fairness” to mean that nobody has any financial loss. However, gains and losses should be the result of investment decisions, rather than unexpected code behaviors.
6.2 Applying TCT on Uniswap
We target the UniswapV2-Solc0.8 implementation [30], which consists of four contracts: the pair contract implements a pair as described above; the UniswapERC20 contract implements the liquidity token for a pair; the factory contract implements the whole market, which creates and stores all pairs; the router contract provides entry functions for the user. Because a liquidity token is specific to a pair, the code derives the pair contract from the UniswapERC20 contract. Thus, an instance of the pair contract behaves both as a reserve of two tradable tokens, and as its own liquidity token.
8 shows the properties defined as postconditions for functions addLiquidity and swap in the router contract. The postcondition of removeLiquidity is identical to that of addLiquidity, so it is not shown. The declaration and assignment lines are used to make the postconditions concise. The operator old is a notation of Boogie. TCT uses it to represent the value of a variable at the beginning of the transaction execution. Note that pair.totalSupply is the amount of liquidity token that this pair contract has minted.
Because these postconditions are defined over three contracts: tokenA, tokenB and their pair (i.e., pair), it is problematic for a static code verifier to examine the safety of the Uniswap’s own contracts (i.e., pair, factory and router) alone, without the concrete code of TokenA and TokenB. For example, a hypothetical TokenA implementation can simply change tokenA.balanceOf[pair] to violate the postconditions, which is the external code totally out of the Uniswap programmers’ control. As we discuss earlier, this is a dilemma for static verification because it needs to make assumptions about external and future implementations.
Our prototype uses a simple token code to construct both TokenA and TokenB. The code is derived from the ERC20 base contract, which specifies the invariants in 6. TCT guarantees that every transaction satisfies the properties of both Uniswap and the tokens.
///@custom:declaration var factory,pair: address;
///@custom:assignment factory:=this.factory; pair:=factory.getPair[tokenA][tokenB];
///@custom:postcondition tokenA.balanceOf[pair]/old(tokenA.balanceOf[pair]) == tokenB.balanceOf[pair]/old(tokenB.balanceOf[pair])
///@custom:postcondition tokenB.balanceOf[pair]/old(tokenB.balanceOf[pair])==pair.totalSupply/old(pair.totalSupply)
function addLiquidity(address tokenA, address tokenB, uint256 amountADesired,
uint256 amountBDesired, uint256 amountAMin,
uint256 amountBMin, address to
///@custom:declaration var factory,pair:address;
///@custom:assignment factory:=this.factory; pair:=factory.getPair[tokenA][tokenB];
///@custom:postcondition old(tokenA.balanceOf[pair]) * old(tokenB.balanceOf[pair]) == tokenA.balanceOf[pair]*tokenB.balanceOf[pair]
function swap(
uint256 amountIn, uint256 amountOutMin,
address tokenA, address tokenB, address to
Code 8: The properties for liquidity and swap operations.
6.3 Higher complexity
Uniswap uses more Solidity language features than those in the studied ERC20 contracts. They include loops, libraries, two-dimensional mappings and dynamic arrays. Every liquidity or swap transaction makes inter-contract calls across many contracts. Our concolic executor correctly converts these complex EVM traces into Boogie programs. Section 7 will explain that the hypotheses of the theorems and the underlying mathematics for the proofs are also much more complex.
7 Evaluation
In this section, we provide quantitative measurements and detailed analyses to show the code complexity that the TCT prototype has handled, the runtime overhead it has achieved, and new code insights it has revealed.
7.1 Code complexity
Table I shows the lines-of-code (LoC) numbers of six EVM traces. The first five rows are the tranfer transactions of the top-5 ERC tokens. Next, TrfrProxy corresponds to the transaction calling MultiVulnToken::transferProxy. Rntrcy and NoRntrcy call MultiVulnToken::clear, the former exploiting the reentrancy pattern and the latter does not. The other traces correspond to Uniswap transactions addLiquidity, removeLiquidity and swap, which are much more complex. For example, Swap contains 9495 EVM instructions involving 5 different contracts. It makes 10 cross-contract calls. The concolic executor generates 299 lines of Boogie code, excluding variable declarations and axioms. The appendix shows the Boogie code for TrfrProxy as an example. All other traces are available in a GitHub repo which we will disclose in the future.
TABLE I: Lines of code of the traces.
Trace
LoC of
EVM code
# of
contracts
# of cross-
contract calls
LoC of
Boogie
USDTTrfr 904 1 0 72
BNBTrfr 686 1 0 64
USDCTrfr 736 1 0 46
stEthTrfr 1119 1 0 60
TONCOINTrfr 683 1 0 32
TrfrProxy 704 1 0 47
Rntrcy 1385 2 3 52
NoRntrcy 628 2 1 21
AddLqdty 7592 5 10 223
RmvLqdty 6985 5 10 199
Swap 9495 5 10 299
7.2 Complexity of the hypotheses and math
Table II shows the hypotheses for the Uniswap traces. They are much more complex than that of TrfrProxy explained in the previous sections. However, with the understanding about Uniswap code, we believe that none of these individual conditions is counterintuitive to the Uniswap programmers. The TCT approach simply explicates their intentions. Under the hypotheses, the traces go through non-trivial computations, equivalent to the math equations shown in the table. Note that tx.origin is a well-defined EVM value, which represents the address of the transaction issuer, i.e., the msg.sender of the transaction’s entry function.
TABLE II: Hypotheses of AddLqdty, RmvLqdty, Swap.
AddLqdty tokenA != tokenB &&
pair.totalSupply > 0 && tx.origin != pair &&
pair.token1 == tokenA && pair.token0 == tokenB &&
tokenA.balanceOf[pair] > 0 && tokenB.balanceOf[pair]
> 0 && tokenA.totalSupply < $\texttt{2}^{\texttt{255}}$& tokenB.totalSupply
< $\texttt{2}^{\texttt{255}}$& pair.reserve0 == tokenB.balanceOf[pair] &&
pair.reserve1 == tokenA.balanceOf[pair]
$\frac{y+a}{y}=\frac{x+a\left(x/y\right)}{x}=\frac{s+{s\left((y+a)-y\right)}/y}{s}$
RmvLqdty tokenA != tokenB &&
tx.origin != pair && pair.totalSupply > 0 &&
to != pair && tokenA.balanceOf[pair] > 0 &&
tokenB.balanceOf[pair] > 0 &&
pair.balanceOf[pair] + liquidity > 0 &&
tokenA.totalSupply<$\texttt{2}^{\texttt{255}}$& tokenB.totalSupply<$\texttt{2}^{\texttt{255}}$
&& pair.reserve0 == tokenB.balanceOf[pair] &&
pair.reserve1 == tokenA.balanceOf[pair] &&
pair.token0 == tokenB && pair.token1 == tokenA
$\frac{x-((b+q)\times x)/s}{x}=\frac{y-((b+q)\times y)/s}{y}=\frac{s-(b+q)}{s}$
Swap factory.swapFeeRate == 0 &&
to != pair && tx.origin != pair &&
pair.token1 == tokenA && pair.token0 == tokenB &&
pair.reserve0 == tokenB.balanceOf[pair] &&
pair.reserve1 == tokenA.balanceOf[pair] &&
tokenB. totalSupply<$\texttt{2}^{\texttt{255}}$& tokenA.totalSupply<$\texttt{2}^{\texttt{255}}$
$x\times y=\left(x-\frac{(a\times 1000)\times x}{(y+a)\times 1000}\right)\times(y+a)$
Each of these hypotheses contains the condition pair.token0 == tokenB && pair.token1 == tokenA. Actually, there are three dual-hypotheses, which Table II does not show due to the space constraint. They contain the condition pair.token0 == tokenA && pair.token1 == tokenB. Hence, each pair contract needs six theorems to cover normal liquidity and swap transactions.
The hypotheses and math reveal the precise relation between the design-level properties and the code. There are two discrepancies to bridge. First, the formulas in Table II contain many divisions, but the code performs integer arithmetic. Concretely checking conditions $x_{2}/x_{1}=y_{2}/y_{1}$ and $x_{1}\times y_{1}=x^{\prime}_{1}\times y^{\prime}_{1}$ in Figure 9 will result in false due to the accumulated imprecision of integer operations. Second, for every swap transaction, a small processing fee is charged by the pair as the incentive for the LPs. The rate is set by factory.swapFeeRate, which is usually set to 3 to specify a 0.3% rate. The Swap code trace is equivalent to the formula in Table II only when the fee rate equals 0 precisely. Therefore, the common wisdom that “Uniswap’s swap transaction obeys the property $x_{1}\times y_{1}=x^{\prime}_{1}\times y^{\prime}_{1}$” is not precise. The actual meaning is the conjunction of two statements: (1) the code trace of the transaction is identical to the code trace of a zero-fee-rate transaction, (2) the latter trace, when assuming the real number arithmetic, implies the precise equality $x_{1}\times y_{1}=x^{\prime}_{1}\times y^{\prime}_{1}$. These statements can only be checked symbolically. In TCT, the path hash mechanism and Boogie’s data type real are used. They bridge the non-trivial gap between the design and the code.
7.3 Runtime overhead
The measurements in Table III and Table IV are made using an HP Z4 G4 Workstation with a 3.80 GHz CPU and 32 GB RAM. Table III shows the runtime overhead on Geth 1.13. The Rntrcy transaction is not included because we intentionally make a false theorem for this case. Also, we do not include the five transfer transactions in the measurements because they are very similar to TrfrProxy. We use the Golang utility time.Now().UnixNano() for the measurement. The baseline is the unmodified Geth. Every measurement is repeated using 20 transactions. Each cell reports the average and the percentage of the standard deviation over the average.
The TCT overhead comes from the hypothesis checking and the path hash computation. For each transaction, these two computations are repeated $10^{6}$ times for higher precision of the measurement. The overhead is the sum of these two numbers divided by the baseline. TrfrProxy and NoRntrcy are simple token transactions. The overheads are both 0.20%. As a comparison, Solythesis [13] is a Solidity code instrumentation technique to insert runtime checks to enforce token properties similar to Prop. 1. It reports 24% slowdown in transaction processing.
TABLE III: Overhead on Geth 1.13 (in nanoseconds).
Trace
Unmodified
Geth (baseline)
Hypothesis
checking
Path hash
computing
Overhead
TrfrProxy $548675\pm 7\%$ $267\pm 16\%$ $839\pm 15\%$ 0.20%
NoRntrcy $608630\pm 11\%$ $185\pm 17\%$ $1005\pm 9\%$ 0.20%
AddLqdty $1517825\pm 17\%$ $4735\pm 7\%$ $3180\pm 8\%$ 0.52%
RmvLqdty $1398780\pm 20\%$ $5242\pm 6\%$ $2832\pm 10\%$ 0.58%
Swap $1289805\pm 18\%$ $3742\pm 6\%$ $3627\pm 7\%$ 0.57%
TABLE IV: Concolic execution time and proof time.
Trace Concolic execution time Boogie proof time
TrfrProxy 3.83 sec 0.94 sec
Rntrcy 3.77 sec 0.80 sec
NoRntrcy 4.07 sec 1.07 sec
AddLqdty 4.42 sec 2.16 sec
RmvLqdty 4.59 sec 5.70 sec
Swap 4.92 sec 1.35 sec
Table IV reports the concolic execution time and the Boogie proof time. Compared to typical formal verification tasks, these numbers are quite low because the reasoning is only about a straight-line code. On the other hand, they are five orders of magnitude higher than those in Table III. The theorem repo turns them into a one-time cost rather than a per-transaction cost. For normal transactions that the programmers anticipate, this one-time cost is paid at the testing time before the contracts process real user transactions.
7.4 Number of theorems expected
The transactions entering from transferProxy of MultiVulnToken and from transfer of BNB, USDC, stETH and TONCOIN require only one theorem, as there is no branch. Function transfer of USDT has a branch, so it needs two theorems (actually, one theorem with two path hashes). Each Uniswap transaction entering from Swap, AddLiquidity, or RemoveLiquidity needs two dual theorems for each token-pair, as explained earlier. In other words, the total number of theorems needed is six times the total number of the pair contracts in order to cover these three operations. Having a linear relation is important for TCT to be applicable to a given scenario. We show that this is true for the studied cases. In Section 8, we propose future work for maintaining the linear relation between the number of theorems and the number of contracts for more complex potential scenarios.
7.5 Insightful violation conditions
TCT’s process explicates insightful conditions, which are often below the design level, such as tx.origin != pair and to != pair for Swap. The former states that the swap transaction must not be issued by the pair itself, and the latter states that the account receiving the swapped-out token from the pair must not be the pair itself. When the conditions are violated, the inverse proportionality will not hold. To show this, we issue six swap transactions to Geth. The reserves of the pair (i.e., $x$ and $y$) are shown in Table V. The former three transactions conform to to != pair, and the latter three violate it. The deviation is the current $x\times y$ divided by the previous $x\times y$ then subtracting 1. Inverse proportionality implies zero deviation. Because of the 0.3% fee, the deviations are [0.0008 , 0.0010] in the second and third rows. However, with to = pair for the last three rows, the value of $y$ becomes unchanged, because the amount of the swapped-out token goes back to the pair. This makes the deviation very large (in fact, arbitrarily large). Hence, if an investment strategy depends on Uniswap unconditionally ensuring inverse proportionality, it can be dangerous. This is similar to integer overflow and reentrancy, as the condition is a detail not foreseeable at the design level. With TCT, specifying inverse proportionality by the designer leads to the explication of conditions like this.
TABLE V: Inverse proportionality violated when to=pair.
$x$ $y$ $x\times y$ Deviation
2000 1000 2000000 N/A
2010 996 2001960 0.0010
2030 987 2003610 0.0008
2050 987 2023350 0.0099
2070 987 2043090 0.0098
2190 987 2161530 0.0580
8 Roadmap Ahead
TCT is a long-term vision for the community. Our prototype shows that it points to a promising direction. Two sub-problems need significant emphasis when we deploy TCT in the real-world setting.
8.1 Consensus of theorem-proving
The Ethereum nodes need to have a consensus on which theorems have been proven. Boogie’s underlying SMT prover is Z3 [31], which is deterministic (unless timeout-based proving tactics are enabled). Similar to Ethereum’s gas mechanism for transaction execution, a prover-gas mechanism should be designed so that all Z3 provers will yield the same result: “sat”, “unsat” or “out of prover gas” for a given SMT problem. Note that some slow nodes falling behind the pace of the majority is a normal situation in Ethereum’s operation, which does not prevent the consensus to be formed. Another potential mechanism to consider is to create a dedicated prover-node consortium, which forms the consensus about which theorems have been proven. Normal Ethereum nodes no longer need to run Z3. Their theorem repositories only need to sync with the latest published snapshot of the consortium’s repository.
8.2 Carrying multiple theorems
In the scenarios that we have discussed, the number of theorems and the number of contracts have a nice linear relation. However, we can imagine scenarios with an exponential number of paths corresponding to independent program branches. We introduce two concepts: “summary” and “carrying multiple theorems”.
Branches often cause local divergence. These so-called diamond splitting can be eliminated by “summaries”. For example, the splitting “if (b) a=1; else a=2;” can be summarized as “modifies(a); assume(a==1 || a==2);”, meaning that no variable other than a is modified, and a is set to either 1 or 2. Reference [32] introduces the notion of summaries for symbolic execution.
Currently, every transaction only carries one theorem. It can be extended to multiple theorems for a transaction shown in Figure 10. ContractA::FuncA calls three cross-contract functions sequentially. The transaction can carry theorem2, theorem3 and theorem4 to prove the postconditions of the three functions. Then, these postconditions are used as the summaries of the functions when proving theorem1 about the safety of running ContractA::FuncA. With this mechanism, the whole transaction is broken into four traces, so the combinatorial complexity is avoided.
Figure 10: A transaction to be proven by multiple theorems.
9 Related Work
TCT is inspired by PCC, PCA and concolic testing. In addition, we summarize the following related technologies.
Static verification for smart contracts. Oyente [33] and Mythril [34] are symbolic verifiers for EVM bytecode. Solidity SMT Checker [35] is a feature built into the Solidity compiler to check assertions and alert possibilities of reentrant calls. Halmos [36] is a recent symbolic analyzer for EVM bytecode compiled from Solidity. In our prototype, we consulted the Halmos code for the EVM semantics.
VerX [37] is a static verifier to prove temporal properties for a set of contracts, with certain assumptions about the “external” contracts that they interact with. The temporal logic is focused on properties that correlate contract states across different transactions. It can reason about liveness properties, which TCT is unable to. Annotary [38] is a Mythril-based verifier that can examine cross-contract transactions. To tackle the state explosion problem, Annotary uses different strategies to treat different variables as either concrete or symbolic, depending on their data types. Note that the authors refer to this approach as “concolic execution”. The meaning is very different from ours.
As explained in the introduction, these static verification techniques consider one or a few “contracts of interest” as the program. Without the runtime info of the contract reference topology, any function making a cross-contract call has arbitrary semantics because the reference is essentially dangling, so some assumptions have to be made about the unknown. Moreover, modeling hypothetical malicious users is a job on thin ice, which leads to false positives and false negatives. These techniques provide offline advisories to programmers, unlike TCT’s runtime certification goal.
Code instrumentation. The Solythesis approach [13] is mentioned earlier. It instruments Solidity code to insert “delta updates” so that aggregation properties, such as sum-of-map, count-of-subset, can be calculated with much lower cost than naively computing the aggregations. The idea is to add auxiliary arrays and maps to keep track of incremental updates of the aggregation quantities when the original code updates the corresponding array/map elements. It is unclear whether Solythesis handles multi-contract situations, as the evaluation is only conducted on 23 individual token contracts. The instrumented contracts show a 24% slowdown.
Other related techniques. The essence of concolic approaches is to take advantage of the logic equivalence between different concrete executions. Some techniques share this essence. Forerunner [39] utilizes the logic equivalence of code traces and data dependencies to enable Ethereum to perform speculative execution for transaction processing. CST [40] is a technique to prove global properties for multi-party web protocol programs, such as OAuth and single-sign-on code. Every a protocol execution reassembles the sequence of the function-calls in the involved parties. The sequence is symbolically examined against the multi-party properties to prove its security or expose bugs.
10 Conclusion
Trustworthiness of the smart contract paradigm requires the community to shift focus from code-level intricacies to design-level properties. The key enabler of the shift will be a solution to the problem that we formulate as “ensuring GP’s conformance to P”, where $GP$ is the self-evolving polymorphic OO program co-developed by the community, and $P$ is the big set of properties co-defined by the same.
This paper shows that TCT is a promising pathway toward the vision. We first demonstrate it on various ERC20 tokens. By holding the transactions against intuitive design-level properties, attacks exploiting unexpected code-level anti-patterns are thwarted. We then successfully apply TCT to highly complex contract code like Uniswap-V2, exemplifying a scale of $\sim$10,000 EVM instructions per transaction across multiple contracts. The near-zero runtime overhead is a crucial contributor to TCT’s practicality. With the TCT protocol in the runtime platform, everyone can feel confident that the code-level computations conform flawlessly to the contracts’ algorithmic designs.
Appendix A Appendix
This appendix provides the automatically-generated Boogie code from the “TrfrProxy” trace in Table I. It helps the understanding about Section 5 and Section 7 of the paper. The code is shown in its entirety. The “LoC of Boogie” number shown in Table I is 47, which only counts the lines after all variable declarations.
type uint256 = int;
const Zero : uint256;
axiom Zero == 0;
const TwoE8 : uint256;
axiom TwoE8 == 32768;
type address = int;
type bytes32 = int;
type uint8 = int;
const TwoE16 : uint256;
axiom TwoE16 == TwoE8 * TwoE8;
const TwoE64 : uint256;
axiom TwoE64 == TwoE16 * TwoE16 * TwoE16 * TwoE16;
const TwoE255 : uint256;
axiom TwoE255 == TwoE64 * TwoE64 * TwoE64 * TwoE16 * TwoE16 * TwoE16 * TwoE8;
const TwoE256 : uint256;
axiom TwoE256 == TwoE64 * TwoE64 * TwoE64 * TwoE64;
function evmadd(a,b:uint256) returns (uint256);
axiom (forall a,b: uint256 :: a+b < TwoE256 && a+b>=Zero ==> evmadd(a,b) == a+b);
axiom (forall a,b: uint256 :: a+b >= TwoE256 && a+b>=Zero ==> evmadd(a,b) == a+b-TwoE256);
axiom (forall a,b: uint256 :: evmadd(a,b)>=a ==> evmadd(a,b) == a+b);
function evmsub(a,b:uint256) returns (uint256);
axiom (forall a,b: uint256 :: a-b < TwoE256 && a>=b ==> evmsub(a,b) == a-b);
axiom (forall a,b: uint256 :: a-b < TwoE256 && a<b ==> evmsub(a,b) == a-b+TwoE256);
axiom (forall a,b: uint256 :: evmsub(a,b)<=a ==> evmsub(a,b) == a-b);
function evmmod(a,b:uint256) returns (uint256);
function sum(m: [address] uint256) returns (uint256);
axiom (forall m: [address] uint256, a:address, v:uint256 :: sum(m[a:=v]) == sum(m) - m[a] + v);
axiom (forall m: [address] uint256 :: ((forall a:address :: Zero<=m[a]) ==> (forall a:address :: m[a]<=sum(m))));
function nondet() returns (uint256);
var Demo.attacker1Address: [address] address;
var Demo.attacker2Address2: [address] address;
var Demo.benignUserAddress2: [address] address;
var MultiVulnToken.owner: [address] address;
var MultiVulnToken.totalSupply: [address] uint256;
var MultiVulnToken.balances: [address] [address] uint256;
var StandardToken.owner: [address] address;
var StandardToken.totalSupply: [address] uint256;
var StandardToken.balances: [address] [address] uint256;
var Token.owner: [address] address;
var Token.totalSupply: [address] uint256;
var no_reentrancy_attack._to: [address] address;
var reentrancy_attack._to: [address] address;
var reentrancy_attack.count: [address] uint256;
var _from: address;
var _to: address;
var _value: uint256;
var _fee: uint256;
procedure straightline_code ()
modifies Demo.attacker1Address, Demo.attacker2Address2, Demo.benignUserAddress2, MultiVulnToken.owner, MultiVulnToken.totalSupply, MultiVulnToken.balances, StandardToken.owner, StandardToken.totalSupply, StandardToken.balances, Token.owner, Token.totalSupply, no_reentrancy_attack._to, reentrancy_attack._to, reentrancy_attack.count, _from, _to, _value, _fee;
var tx_origin: address;
var entry_contract: address;
var BLOCKTIME: uint256;
var tmp1: uint256;
var tmp2: uint256;
var tmp3: bool;
var tmp4: bool;
var tmp5: uint256;
var tmp6: uint256;
var tmp7: uint256;
var tmp8: bool;
var tmp9: bool;
var tmp10: uint256;
var tmp11: uint256;
var tmp12: uint256;
var tmp13: bool;
var tmp14: bool;
var tmp15: uint256;
var tmp16: uint256;
var tmp17: uint256;
var tmp18: uint256;
var tmp19: uint256;
var tmp20: uint256;
var tmp21: uint256;
// def-vars
var totalSupply: uint256;
totalSupply:= MultiVulnToken.totalSupply[entry_contract];
// hypothesis
assume(Zero <= _value);
assume(_value < TwoE255);
assume(Zero <= _fee);
assume(_fee < TwoE255);
assume(totalSupply < TwoE255);
// insert invariant of entry contract
assume(forall x:address :: Zero <= MultiVulnToken.balances[entry_contract][x] && MultiVulnToken.balances[entry_contract][x] <= MultiVulnToken.totalSupply[entry_contract]);
assume(sum( MultiVulnToken.balances[entry_contract] ) == MultiVulnToken.totalSupply[entry_contract]);
tmp1:=MultiVulnToken.balances[entry_contract][_from];
tmp2:=evmadd(_fee,_value);
tmp3:= (tmp1<tmp2);
tmp4:=!tmp3;
assume(tmp4);
tmp5:=MultiVulnToken.balances[entry_contract][_to];
tmp6:=MultiVulnToken.balances[entry_contract][_to];
tmp7:=evmadd(tmp6,_value);
tmp8:= (tmp7<tmp5);
tmp9:=!tmp8;
assume(tmp9);
tmp10:=MultiVulnToken.balances[entry_contract][tx_origin];
tmp11:=MultiVulnToken.balances[entry_contract][tx_origin];
tmp12:=evmadd(tmp11,_fee);
tmp13:= (tmp12<tmp10);
tmp14:=!tmp13;
assume(tmp14);
tmp15:=MultiVulnToken.balances[entry_contract][_to];
tmp16:=evmadd(tmp15,_value);
MultiVulnToken.balances[entry_contract][_to]:=tmp16;
tmp17:=MultiVulnToken.balances[entry_contract][tx_origin];
tmp18:=evmadd(tmp17,_fee);
MultiVulnToken.balances[entry_contract][tx_origin]:=tmp18;
tmp19:=MultiVulnToken.balances[entry_contract][_from];
tmp20:=evmadd(_value,_fee);
tmp21:=evmsub(tmp19,tmp20);
MultiVulnToken.balances[entry_contract][_from]:=tmp21;
// (post) insert invariant of entry contract
assert(forall x:address :: Zero <= MultiVulnToken.balances[entry_contract][x] && MultiVulnToken.balances[entry_contract][x] <= MultiVulnToken.totalSupply[entry_contract]);
assert(sum( MultiVulnToken.balances[entry_contract] ) == MultiVulnToken.totalSupply[entry_contract]);